OpenStudy (anonymous):

Use the quadratic formula to solve the quadratic equation: 2x2 − 3x − 7 = 0

OpenStudy (saifoo.khan):

im getting a decimal answer!

OpenStudy (saifoo.khan):

x = 2.7655 x = -1.2655

OpenStudy (anonymous):

u should \[(3\pm \sqrt{65})/4\]

OpenStudy (anonymous):

Polynomial Equation: \[ax ^{2} + bx + c = 0\] Quadratic equation: \[\left( -b \pm \sqrt{b ^{2}-4ac} \right)\div \left( 2a \right)\] So: a = 2 b = -3 c = -7 That's how jahtoday got the formula above.